Prove that \sum_{i=0}^{n} 2^{i} = 2^{n + 1} - 1 Use mathematical induction for this proof and discuss/explain each step.


1
Expert's answer
2022-02-21T16:42:32-0500

Let us prove that "\sum\limits_{i=0}^{n} 2^{i} = 2^{n + 1} - 1" using mathematical induction.


For "n=1," we have that the left side is "2^0+2^1=3," and the right side is "2^2-1=3." Therefore, for "n=1" the equality is true.


Suppose that the equality is true for "n=k," that is "\sum\limits_{i=0}^{k} 2^{i} = 2^{k + 1} - 1."

Let us prove for "n=k+1."


It follows that


"\sum\limits_{i=0}^{k+1} 2^{i} =\sum\limits_{i=0}^{k} 2^{i}+2^{k+1} = 2^{k + 1}-1+ 2^{k + 1} =2\cdot 2^{k + 1} - 1=2^{k+2}-1."


We conclude that by principle of mathematical induction the statement is true for all natural numbers "n."